OpenStudy (anonymous):

the "table" shows the realtionship between x and y. X: 0 1 2 3 4 5 Y: 1 2 5 10 ? ? what r the next 2 variables? a) 11 and 13 b) 15 and 22 c) 16 and 25 d) 17 and 26 how do you know?

OpenStudy (amistre64):

this is pretty much a sequence generator

OpenStudy (amistre64):

each new y is twice its older partner

OpenStudy (amistre64):

that was a bad read lol

OpenStudy (anonymous):

wha? can u show me like, an equation for it or something?

OpenStudy (amistre64):

10/5 = 2 5/2 = 2.5 2/1 = 2 might be the pattern

OpenStudy (anonymous):

5 is not twice 2

OpenStudy (amistre64):

5/4 might - 2?

OpenStudy (amistre64):

might = 2

OpenStudy (anonymous):

oh! i know how ill solve it! GRAPHING!

OpenStudy (anonymous):

Y=X^2+1 I think so should be d

OpenStudy (amistre64):

jethro saw it ;)

OpenStudy (anonymous):

Y=X^2+1 I think so should be d

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

thanks jethrokaun!